Sol Summary – December 1st

Crew 183+ Sol Summary Report 1 December, 2017

Sol 4

Title: Curtis Arrives!

Written by: Brandon Ferguson

Mission Status: Good. Operations are Normal

Sol Activity Summary:

Today, we spent most of our time at the hab. At long last, Curtis arrives to fix the water heater in the hab, which has been out of commission for over a month now. He also insulated the water pipes and the outside tank. This morning, I was able to start my final batch of Sedimentation tests. All that’s left are my salinity & pH tests for all the soils.

Meanwhile, Chris offered to help put the rovers together. We now have three out of the four new rovers fully assembled.

At about 4:00 PM, I took a short EVA to Tank wash to resupply on a couple of samples that I needed.

Look Ahead Plan:

Tomorrow morning, Curtis will deliver and install Shannon’s water tank tomorrow, Shannon and I will be preparing the Lab and the Hab for Crew 184’s arrival

Anomalies in work: No Problems

Weather: Clear, warm, calm winds

Crew Physical Status: OK, Healthy

EVA: Tank Wash area

Reports to be Filed:

-Operations Report

Support Requested: N/A

Operations Report – December 1st

Crew 183 + Operations Report 1 Dec 2017

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: Phone is still not functional, Low water pressure in Hab

Notes on non-nominal systems: HudgesNet was here all day but unable
to repair the phone. They will be working on it again on Monday. The
filter than normally clogs up the water line in the Hab is missing,
but two other filters were cleaned (Bathroom and kitchen sink) and
water pressure is once again nominal throughout the Hab. A new master
filter will be installed to forever end this problem, which was
probably why the old water heater died.

Generator (hours run): off 8 am on 6 pm

Solar— SOC 94% (Before generator is run at night)

Diesel – 50 %

Propane – 100 psi

Ethanol Free Gasoline (5 Gallon containers for ATV) – not known gallons.

Water (trailer) – 0 gallons. Curtis took the water trailer to town to fill it.

Water (static) – not known gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: not known

Toilet tank emptied: no

ATV’s Used: yes

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 2

Notes on ATVs: nothing to report

Deimos rover used: yes


Beginning charge: 100

Ending charge: 100

Currently charging: no

Sojourner rover used: ASSIGNED TO DIRECTOR, Operational


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Operational


Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: Operational


Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: Not operational


Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where? No

General notes and comments: Remember that for the first month of use,
the new rovers will not hold a charge as well as Deimos. We took the
three new ones on a 5 mile round trip and they were each reduced to
about 75%. It takes a while for the batteries to perform at their

I’d like to say that the rovers were made operational by Brandon and I
but it was actually Chris who did a majority of the work and taught me
a lot about the right tools for the job. Who is Chris? Well, he’s
the guy who was here to fix my phone. It took so long he offered to

The static tank was insulated today. A new heater was installed and
the pipe to the Hab was insulated as well. Shannon’s water tank was
delivered and will be installed tomorrow.

Water heater: After determining that the new water heater would not
fit in the spot where the old one was mounted, it was moved to a spot
on the lower deck just to the right of the ductless propane heater.
It was operational, but would not function correctly until the water
pressure was restored. It is currently connected to the heater’s
propane line. I’m assuming Curtis will reconnect the heater as soon
as he has the supplies.

Summary of internet: not known

Summary of suits and radios: nothing to report

Summary of Hab operations: Water system insulated on the Hab’s
exterior. New hot water heater installed in new location. Water
pressure restored. Water trailer to be refilled.

Summary of GreenHab operations: The thinned plants look mostly good.
One of the tomatoes I tried to transplant does not look good. The
rest look fine. Door was open for a few hours today but it was cold
enough that it stayed mostly warm and not hot.

Summary of ScienceDome operations: Brandon concluded his work.
Nothing more to report.

Summary of RAMM operations: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: Nothing, thank you.

Copyright © The Mars Society. All rights reserved. | Main Site